# Weak Convexity of Fisher Information Matrix and Superresolved Localization of Blinking Sources of Light

### v2

@@ V Localization of two incoherent sources
In this section, we consider a simple example illustrating the theory developed above and consisting of a one-dimensional localization of two incoherent sources lying at unknown locations $\bar{x}_{1}$ and $\bar{x}_{2}$ on the object plane of the microscope and having known relative brightnesses $\mu_{1}$ and $\mu_{2}$. On its image plane, they are represented by two spots centered at $x_{1}=g\bar{x}_{1}$ and $x_{2}=g\bar{x}_{2}$, where $g$ is the microscope magnification. We consider the task of estimating the parameter vector $\theta=\{x_{1},x_{2}\}$, from which $\bar{x}_{1}$ and $\bar{x}_{2}$ can easily be found.
@@
[…]
$a^{\dagger}(x)$ being the photon creation operator at point $x$. In direct detection, the POVM for detecting a photon at position $x$ is the projector on the one-photon state at position $x$, i.e., $\Lambda(x)=a^{\dagger}(x)|\text{vac}\rangle\langle\text{vac}|a(x)$. Then, the probability of measurement outcome $x$ is $p_{k}(x|x_{1},x_{2})=\operatorname{Tr}\left[\rho_{Pk}\Lambda(x)\right]=\psi^{2}(x-x_{k})$, where we have used the commutation relation $[a(x),a^{\dagger}(x^{\prime})]=\delta(x-x^{\prime})$. The corresponding Fisher information matrix is $F^{[k]}_{ij}=\sigma^{-2}\delta_{ij}\delta_{ik}$. The Fisher information matrix of the entire estimation session with $N$ detected photons, $\mu_{1}N$ of which coming from the first source and $\mu_{2}N$ coming from the second one, is thus
$\mathbf{F}_{\text{blink}}=\frac{N}{2\sigma^{2}}\left(\begin{array}[]{cc}1+\delta&0\\ 0&1-\delta\end{array}\right),$ (45)
where we express the relative brightnesses, satisfying the relation $\mu_{1}+\mu_{2}=1$, through one parameter $10<\delta<$ 1, which we assume to be known, as $\mu_{1,2}=\frac{1}{2}(1\pm\delta)$. The average Fisher information in this scenario is $\bar{F}_{\text{blink}}=N/2\sigma^{2}$.
@@
[…]
Figure 3: Variance lower bounds in the units of $L_{0}=2\sigma^{2}/N$ for estimating the unknown positions $(x_{1},x_{2})$ of two incoherent sources by detecting $N$ photons in a microscope with the point-spread function of standard deviation $\sigma$. $\delta$ is the relative brightness difference of the sources. Parameters $(c,s)$ are the scaled centroid and separation, respectively. Parameters $(b,w)$ are the eigenparameters of the Fisher information matrix obtained from $(x_{1},x_{2})$ via a rotation by angle $\xi$.
The corresponding variance lower bounds $L_{ii}=(\mathbf{F}^{-1})_{ii}$ are calculated numerically and shown in Fig. 3(a) for the case of equal brightnesses $\mu_{1}=\mu_{2}=\frac{1}{2}$. In this case, due to the symmetry with respect to the exchange of parameters $x_{1}$ and $x_{2}$, we have $L_{11}=L_{22}$ for both scenarios, and the variance of the cofluorescence scenario is always above that of the blinking scenario, in correspondence with Theorem 2.
@@
We also see in Fig. 3(a) that the variances of both positions tend to infinity at a small separation, which means that these positions are difficult to estimate. However, intuitively we understand that the centroid $(x_{1}+x_{2})/2$ of the overlapping spots can still be estimated with high precision. To show it explicitly, we pass from our parameters $\theta=\{x_{1},x_{2}\}$ to a rotated vector of parameters $\varphi=\{c,s\}$, where $c=(x_{1}+x_{2})/\sqrt{2}$ is the scaled centroid and $s=(x_{2}-x_{1})/\sqrt{2}$ is the scaled separation. This rotation in the parameter space can be written as $\varphi=O_{\pi/4}^{T}\theta$, where
$O_{\xi}=\left(\begin{array}[]{cc}\cos\xi&-\sin\xi\\ \sin\xi&\cos\xi\end{array}\right)$ (48)
is the matrix of orthogonal rotation by angle $\xi$. The Fisher information matrix in the new basis is simply $\tilde{\mathbf{F}}=O_{\pi/4}^{T}\mathbf{F}O_{\pi/4}$. Applying this transformation to Eq. (45), we obtain
$\tilde{\mathbf{F}}_{\text{blink}}=\frac{N}{2\sigma^{2}}\left(\begin{array}[]{cc}1&-\delta\\ -\delta&1\end{array}\right),$ (49)
@@
$\lim_{x_{2}\to x_{1}}\mathbf{F}_{\text{coflu}}=\frac{N}{\sigma^{2}}\left(\begin{array}[]{cc}\mu_{1}^{2}&\mu_{1}\mu_{2}\\ \mu_{1}\mu_{2}&\mu_{2}^{2}\end{array}\right).$ (50)
This matrix has a zero determinant, but its trace is non-zero. It means that one of its eigenvalues is zero, but another is not: $\lambda_{1}=N(\mu_{1}^{2}+\mu_{2}^{2})/\sigma^{2}$, $\lambda_{2}=0$. As a consequence, at any $s$, we can introduce two eigenparameters, as discussed in Sec. II.2: the best combination of the two positions $b=x_{1}\cos\xi+x_{2}\sin\xi$ corresponding to the higher eigenvalue of $\mathbf{F}_{\text{coflu}}(s)$ and the weighted separation $w=x_{2}\cos\xi-x_{1}\sin\xi$ corresponding to its lower eigenvalue. The latter parameter has a divergent variance at $s\to 0$, while the former is well determined.
In Fig. 3(e), we illustrate the behavior of eigenparameter variances and the angle $\xi$ in the case of equal brightnesses. We see that $\xi=\pi/4$ at low separation, which corresponds to $b=c$ and $w=s$, exactly as in Fig. 3(c). At higher separation, the angle becomes $\xi=-\pi/4$, which means that the parameters are relabeled: $b=-s$ and $w=c$. This happens because $s$ has a higher eigenvalue in this region. We also see that both eigenvalues of $\mathbf{F}_{\text{coflu}}^{-1}(\theta)$ are larger than the degenerate eigenvalue of $\mathbf{F}_{\text{blink}}^{-1}(\theta)$, as prescribed by Eq. (37).
A more interesting case of unequal brightnesses is illustrated in Fig. 3(f). We see that the angle $\xi$ changes continuously, tending to zero at high separation, where the matrix $\mathbf{F}_{\text{coflu}}(s)$ becomes diagonal in the $(x_{1},x_{2})$ basis. The parameter $b$ has always a finite variance, while that of $w$ is divergent at a small separation. Both variances in the cofluorescent scenario lie above their counterparts in the blinking scenario, which means that $\lambda_{i}(\mathbf{F}_{\text{blink}})\geq\lambda_{i}(\mathbf{F}_{\text{coflu}})$, where $\lambda_{i}(A)$ is the $i$th eigenvalue of a Hermitian matrix $A$, when the eigenvalues are sorted in a nonincreasing order. It is known that for any two Hermitian matrices $A$ and $B$, $A\geq B$ is a sufficient but not necessary condition for $\lambda_{i}(A)\geq\lambda_{i}(B)$ (see Corollary 7.7.4. in Ref. [35]). This means that we cannot conclude from Fig. 3(f) that inequality (37) holds also in the case of unequal brightnesses, not comprised by Theorem 2. However, an additional numerical study shows that the minimal eigenvalue of the matrix $\mathbf{F}_{\text{blink}}-\mathbf{F}_{\text{coflu}}$ is positive in the limits of computational error, which is a numerical evidence in favor of validity of inequality (37) in a more general case of unequal brightnesses.
@@ VI.1 Additivity and convexity of quantum Fisher information matrix
Given a positive integer $M$, a state space of a quantum system $\mathcal{S}$, and a family of density operators $\rho(\theta)\in\mathcal{S}$ depending on parameter $\theta\in\mathbb{R}^{M}$, the quantum Fisher information matrix for the information contained in $\rho(\theta)$ about $\theta$ is the $M\times M$-matrix-valued map $\mathcal{Q}_{M}^{(\mathcal{S})}\left[\rho(\theta)\right]$, whose matrix elements are
@@
$\mathcal{Q}_{M}^{(\mathcal{S}_{1}\otimes\mathcal{S}_{2})}\left[\rho_{1}(\theta)\otimes\rho_{2}(\theta)\right]$ (53)
$=\mathcal{Q}_{M}^{(\mathcal{S}_{1})}\left[\rho_{1}(\theta)\right]+\mathcal{Q}_{M}^{(\mathcal{S}_{2})}\left[\rho_{2}(\theta)\right],$
@@
for any $10<\gamma<$ 1 and $\rho_{1}(\theta)$ and $\rho_{2}(\theta)$ satisfying the requirement
@@
Given two positive integers $K\geq 2$ and $d$, in the general case of $M$ parameters split into $K$ groups of $d$ parameters each, $\theta=\{\bar{\theta}_{1},...,\bar{\theta}_{K}\}$, where $\bar{\theta}_{k}\in\mathbb{R}^{d}$, $M=Kd$, and given real numbers $\mu_{1}$,…,$\mu_{K}$ belonging to $[0,1)$ and summing up to unity, the quantum Fisher information matrix possesses the following properties.
@@
    $\mathcal{Q}_{M}^{(\mathcal{S})}\left[\frac{1}{K}\sum\limits_{k=1}^{K}\rho_{k}(\theta)\right]<\frac{1}{K}\sum\limits_{k=1}^{K}\mathcal{Q}_{M}^{(\mathcal{S})}\left[\rho_{k}(\theta)\right],$ (59)
@@
    $\operatorname{Tr}\left\{\mathcal{Q}_{M}^{(\mathcal{S})}\left[\sum\limits_{k=1}^{K}\mu_{k}\rho_{k}(\theta)\right]\right\}<\sum\limits_{k=1}^{K}\mu_{k}\operatorname{Tr}\left\{\mathcal{Q}_{M}^{(\mathcal{S})}\left[\rho_{k}(\theta)\right]\right\}.$ (61)
@@ VI.2 Quantum Fisher information for localization of blinking sources
The consideration of Sec. IV can be extended to quantum Fisher information. In the cofluorescence scenario, the state of the probe, defined in Sec. II.1, is $\tilde{\bm{\rho}}_{P}^{\text{coflu}}=\left(\tilde{\rho}_{P}^{\text{coflu}}\right)^{\otimes N}$, where
@@
and the quantum Fisher information matrix, after Lemma 4, has the form $\mathbf{Q}_{\text{coflu}}(\theta)=NQ_{\text{coflu}}(\theta)$, where $Q_{\text{coflu}}(\theta)$ is the single-sample quantum Fisher information matrix with the elements
@@
$\tilde{\bm{\rho}}_{P}^{\text{blink}}=\tilde{\rho}_{P1}^{\otimes N_{1}}\otimes...\otimes\tilde{\rho}_{PK}^{\otimes N_{K}}$ (64)
and the quantum Fisher information matrix is $\mathbf{Q}_{\text{blink}}(\theta)=\sum_{k}N_{k}Q^{[k]}(\theta)$, where $Q^{[k]}(\theta)$ is the single-sample quantum Fisher information matrix in the $k$th window, where just the $k$th source is active:
@@
For $\mathbf{Q}_{\text{coflu}}(\theta)$ defined by Eq. (14) with the object state (62) and $\mathbf{Q}_{\text{blink}}(\theta)$ defined by Eq. (14) with the object state (64), the following inequalities hold:
@@
    $\mathbf{Q}_{\text{blink}}(\theta)>\mathbf{Q}_{\text{coflu}}(\theta).$ (66)
@@
    $\operatorname{Tr}\{\mathbf{Q}_{\text{blink}}(\theta)\}>\operatorname{Tr}\{\mathbf{Q}_{\text{coflu}}(\theta)\}.$ (67)
@@
As a corollary to Theorem 4, we have $\mathbf{Q}_{\text{blink}}^{-1}(\theta)<\mathbf{Q}_{\text{coflu}}^{-1}(\theta)$ in case A. Does it mean that the localization efficiency is always higher in the blinking scenario for any choice of measurement? In the case of Fisher information discussed in Sec. IV, the answer was positive, because the Cramér-Rao bound is always asymptotically saturable. The quantum Cramér-Rao bound, however, is asymptotically saturable only in the case where the equality
$\operatorname{Tr}\left\{\left(\mathcal{L}_{i}\mathcal{L}_{j}-\mathcal{L}_{j}\mathcal{L}_{i}\right)\tilde{\bm{\rho}}_{P}(\theta)\right\}=0$ (68)
@@
[…]
@@
where $i$ takes the values $x_{k},y_{k}$ and the states $|\partial_{x}\Psi_{k}\rangle$ and $|\partial_{y}\Psi_{k}\rangle$ are defined similarly to Eq. (69) but with replacements $\psi(x,y)\to\partial_{x}\psi(x,y)$ and $\psi(x,y)\to\partial_{y}\psi(x,y)$, respectively. Note, that the square of the function $\psi(x,y)$ is normalized to unity, but the square of its derivative is not necessarily that, i.e., the state $|\partial_{i}\Psi_{k}\rangle$ is not normalized in general. Differentiating the norm $\langle\Psi_{k}|\Psi_{k}\rangle=1$, we obtain $\operatorname{Re}\left\{\langle\partial_{i}\Psi_{k}|\Psi_{k}\rangle\right\}=0$, and since the impulse-response function is assumed to be real, $|\partial_{i}\Psi_{k}\rangle$ is orthogonal to $|\Psi_{k}\rangle$. Substituting Eq. (70) into Eq. (68), we ascertain that the commutator average is zero. It means that $x_{k}$ and $y_{k}$ are compatible for a simultaneous quantum estimation in the window where they are active. The same argument applies to the locations of all other sources.
In summary, the quantum Cramér-Rao bound is asymptotically saturable in the blinking scenario. We do not discuss here whether it is saturable or not in the cofluorescence scenario, because, in any case, the advantage of the blinking scenario, expressed by inequality (42), holds for the optimal quantum measurement. As in Sec. IV, this inequality holds for any measure of localization efficiency in the case of equal brightnesses and a quantum translation and rotation invariant microscopy, and for $H_{\text{eig}}(\hat{\theta})$ in the general case.
It is easy to see that the property of quantum translation and rotation invariance, formulated in Definition 7, is reached for a rotationally invariant impulse-response function, $\psi(Rr)=\psi(r)$, where $R$ is a matrix of rotation on the $(x,y)$ plane. Indeed, from Eqs. (69) and (70), we obtain $\mathcal{Q}_{2}^{(\mathcal{S})}\left[\tilde{\rho}_{Pk}\right]_{ij}=J_{ij}$, where

### v3

@@ V Localization of two incoherent sources
In this section, we consider a simple example illustrating the theory developed above and consisting of a one-dimensional localization of two incoherent sources lying at unknown locations $\bar{x}_{1}$ and $\bar{x}_{2}$ on the object plane of the microscope and having known relative brightnesses $\mu_{1}$ and $\mu_{2}$. On its image plane, they are represented by two spots centered at $x_{1}=g\bar{x}_{1}$ and $x_{2}=g\bar{x}_{2}$, where $g$ is the microscope magnification. We consider the task of estimating the parameter vector $\theta=\{x_{1},x_{2}\}$, from which $\bar{x}_{1}$ and $\bar{x}_{2}$ can easily be found, see Fig. 1.
@@
$|\Psi(x_{k})\rangle=\int\limits_{-\infty}^{+\infty}dx\psi(x-x_{k})a^{\dagger}(x)|\text{vac}\rangle,$ (44)
$a^{\dagger}(x)$ being the photon creation operator at point $x$. In direct detection, the POVM for detecting a photon at position $x$ is the projector on the one-photon state at position $x$, i.e., $\Lambda(x)=a^{\dagger}(x)|\text{vac}\rangle\langle\text{vac}|a(x)$. Then, the probability of measurement outcome $x$ is $p_{k}(x|x_{1},x_{2})=\operatorname{Tr}\left[\rho_{Pk}\Lambda(x)\right]=\psi^{2}(x-x_{k})$, where we have used the commutation relation $a(x),a^{\dagger}(x^{\prime})]=\delta(x-x^{\prime})$. The corresponding Fisher information matrix is $F^{[k]}_{ij}=\sigma^{-2}\delta_{ij}\delta_{ik}$. The Fisher information matrix of the entire estimation session with $N$ detected photons, $\mu_{1}N$ of which coming from the first source and $\mu_{2}N$ coming from the second one, is thus
$\underline{\mathbf{F}}_{\text{blink}}=\frac{N}{2\sigma^{2}}\left(\begin{array}[]{cc}1+\delta&0\\ 0&1-\delta\end{array}\right),$ (45)
where we express the relative brightnesses, satisfying the relation $\mu_{1}+\mu_{2}=1$, through one parameter $0<\delta<1$, which we assume to be known, as $\mu_{1}=\frac{1}{2}(1+\delta)$ and $\mu_{2}=\frac{1}{2}(1-\delta)$. The average Fisher information in this scenario is $\bar{F}_{\text{blink}}=N/2\sigma^{2}$.
@@
$\underline{F}^{\text{coflu}}_{ij}=\frac{N\mu_{i}\mu_{j}}{\sigma^{4}}\int\limits_{-\infty}^{+\infty}(x-x_{i})(x-x_{j})\frac{p_{i}(x|x_{i})p_{j}(x|x_{j})}{p(x|x_{1},x_{2})}dx.$ (47)
Figure 4: Variance lower bounds in the units of $L_{0}=2\sigma^{2}/N$ for estimating the unknown positions $(x_{1},x_{2})$ of two incoherent sources by detecting $N$ photons in a microscope with the point-spread function of standard deviation $\sigma$. $\delta$ is the relative brightness difference of the sources. Parameters $(c,s)$ are the scaled centroid and separation, respectively. Parameters $(b,w)$ are the eigenparameters of the Fisher information matrix obtained from $(x_{1},x_{2})$ via a rotation by angle $\xi$: $b=x_{1}\cos\xi+x_{2}\sin\xi$ and $w=x_{2}\cos\xi-x_{1}\sin\xi$.
The corresponding variance lower bounds $L_{ii}=(\underline{\mathbf{F}}^{-1})_{ii}$ are calculated numerically and shown in Fig. 4(a) for the case of equal brightnesses $\mu_{1}=\mu_{2}=\frac{1}{2}$. In this case, due to the symmetry with respect to the exchange of parameters $x_{1}$ and $x_{2}$, we have $L_{11}=L_{22}$ for both scenarios, and the variance of the cofluorescence scenario is always above that of the blinking scenario, in correspondence with Theorem 2.
@@
We also see in Fig. 4(a) that the variances of both positions tend to infinity at a small separation, which means that these positions are difficult to estimate. However, intuitively we understand that the centroid $(x_{1}+x_{2})/2$ of the overlapping spots can still be estimated with high precision. To show it explicitly, we pass from our parameters $\theta=\{x_{1},x_{2}\}$ to a rotated vector of parameters $\varphi=\{c,s\}$, where $c=(x_{1}+x_{2})/\sqrt{2}$ is the scaled centroid and $s=(x_{2}-x_{1})/\sqrt{2}$ is the scaled separation. This rotation in the parameter space can be written as $\varphi=\mathbf{O}_{\pi/4}^{T}\theta$, where
$\mathbf{O}_{\xi}=\left(\begin{array}[]{cc}\cos\xi&-\sin\xi\\ \sin\xi&\cos\xi\end{array}\right)$ (48)
is the matrix of orthogonal rotation by angle $\xi$. The Fisher information matrix in the new basis is simply $\tilde{\underline{\mathbf{F}}}=\mathbf{O}_{\pi/4}^{T}\underline{\mathbf{F}}\mathbf{O}_{\pi/4}$. Applying this transformation to Eq. (45), we obtain
$\tilde{\underline{\mathbf{F}}}_{\text{blink}}=\frac{N}{2\sigma^{2}}\left(\begin{array}[]{cc}1&-\delta\\ -\delta&1\end{array}\right),$ (49)
@@
$\lim_{x_{2}\to x_{1}}\underline{\mathbf{F}}_{\text{coflu}}=\frac{N}{\sigma^{2}}\left(\begin{array}[]{cc}\mu_{1}^{2}&\mu_{1}\mu_{2}\\ \mu_{1}\mu_{2}&\mu_{2}^{2}\end{array}\right).$ (50)
This matrix has a zero determinant, but its trace is non-zero. It means that one of its eigenvalues is zero, but another is not: $\lambda_{1}=N(\mu_{1}^{2}+\mu_{2}^{2})/\sigma^{2}$, $\lambda_{2}=0$. As a consequence, at any $s$, we can introduce two eigenparameters, as discussed in Sec. II.2: the best combination of the two positions $b=x_{1}\cos\xi+x_{2}\sin\xi$ corresponding to the higher eigenvalue of $\underline{\mathbf{F}}_{\text{coflu}}(s)$ and the weighted separation $w=x_{2}\cos\xi-x_{1}\sin\xi$ corresponding to its lower eigenvalue. The latter parameter has a divergent variance at $s\to 0$, while the former is well determined.
In Fig. 4(e), we illustrate the behavior of eigenparameter variances and the angle $\xi$ in the case of equal brightnesses. We see that $\xi=\pi/4$ at low separation, which corresponds to $b=c$ and $w=s$, exactly as in Fig. 4(c). At higher separation, the angle becomes $\xi=-\pi/4$, which means that the parameters are relabeled: b=sb=-s and $w=c$. This happens because $s$ has a higher eigenvalue in this region. We also see that both eigenvalues of $\underline{\mathbf{F}}_{\text{coflu}}^{-1}(\theta)$ are larger than the degenerate eigenvalue of $\underline{\mathbf{F}}_{\text{blink}}^{-1}(\theta)$, as prescribed by Eq. (37).
A more interesting case of unequal brightnesses is illustrated in Fig. 4(f). We see that the angle $\xi$ changes continuously, tending to zero at high separation, where the matrix $\underline{\mathbf{F}}_{\text{coflu}}(s)$ becomes diagonal in the $(x_{1},x_{2})$ basis. The parameter $b$ has always a finite variance, while that of $w$ is divergent at a small separation. Both variances in the cofluorescent scenario lie above their counterparts in the blinking scenario, which means that $\lambda_{i}(\underline{\mathbf{F}}_{\text{blink}})\geq\lambda_{i}(\underline{\mathbf{F}}_{\text{coflu}})$, where $\lambda_{i}(\mathbf{A})$ is the $i$th eigenvalue of a Hermitian matrix $\mathbf{A}$, when the eigenvalues are sorted in a nonincreasing order. It is known that for any two Hermitian matrices $\mathbf{A}$ and $\mathbf{B}$, $\mathbf{A}\geq\mathbf{B}$ is a sufficient but not necessary condition for $\lambda_{i}(\mathbf{A})\geq\lambda_{i}(\mathbf{B})$ (see Corollary 7.7.4. in Ref. [35]). This means that we cannot conclude from Fig. 4(f) that inequality (37) holds also in the case of unequal brightnesses, not comprised by Theorem 2. However, an additional numerical study shows that the minimal eigenvalue of the matrix $\underline{\mathbf{F}}_{\text{blink}}-\underline{\mathbf{F}}_{\text{coflu}}$ is positive in the limits of computational error, which is a numerical evidence in favor of validity of inequality (37) in a more general case of unequal brightnesses.
@@ VI.1 Additivity and convexity of quantum Fisher information matrix
Given a positive integer $M$, a state space of a quantum system $\mathcal{S}$, and a family of density operators $\rho(\theta)\in\mathcal{S}$ depending on parameter $\theta\in\mathbb{R}^{M}$, the quantum Fisher information matrix for the information contained in $\rho(\theta)$ about $\theta$ is the $M\times M$-matrix-valued map $\bm{\mathcal{Q}}_{M}^{(\mathcal{S})}\left[\rho(\theta)\right]$, whose matrix elements are
@@
$\bm{\mathcal{Q}}_{M}^{(\mathcal{S}_{1}\otimes\mathcal{S}_{2})}\left[\rho_{1}(\theta)\otimes\rho_{2}(\theta)\right]$ (53)
$=\bm{\mathcal{Q}}_{M}^{(\mathcal{S}_{1})}\left[\rho_{1}(\theta)\right]+\bm{\mathcal{Q}}_{M}^{(\mathcal{S}_{2})}\left[\rho_{2}(\theta)\right],$
@@
for any $0<\gamma<1$ and $\rho_{1}(\theta)$ and $\rho_{2}(\theta)$ satisfying the requirement
@@
Given two positive integers $K\geq 2$ and $d$, in the general case of $M$ parameters split into $K$ groups of $d$ parameters each, $\theta=\{\bar{\theta}_{1},...,\bar{\theta}_{K}\}$, where $\bar{\theta}_{k}\in\mathbb{R}^{d}$, $M=Kd$, and given real numbers $\mu_{1}$$\mu_{K}$ belonging to $[0,1)$ and summing up to unity, the quantum Fisher information matrix possesses the following properties.
@@
    $\bm{\mathcal{Q}}_{M}^{(\mathcal{S})}\left[\frac{1}{K}\sum\limits_{k=1}^{K}\rho_{k}(\theta)\right]<\frac{1}{K}\sum\limits_{k=1}^{K}\bm{\mathcal{Q}}_{M}^{(\mathcal{S})}\left[\rho_{k}(\theta)\right],$ (59)
@@
    $\operatorname{Tr}\left\{\bm{\mathcal{Q}}_{M}^{(\mathcal{S})}\left[\sum\limits_{k=1}^{K}\mu_{k}\rho_{k}(\theta)\right]\right\}<\sum\limits_{k=1}^{K}\mu_{k}\operatorname{Tr}\left\{\bm{\mathcal{Q}}_{M}^{(\mathcal{S})}\left[\rho_{k}(\theta)\right]\right\}.$ (61)
@@ VI.2 Quantum Fisher information for localization of blinking sources
The consideration of Sec. IV can be extended to quantum Fisher information. In the cofluorescence scenario, the state of the probe, defined in Sec. II.1, is $\tilde{\underline{\rho}}_{P}^{\text{coflu}}=\left(\tilde{\rho}_{P}^{\text{coflu}}\right)^{\otimes N}$, where
@@
and the quantum Fisher information matrix, after Lemma 4, has the form $\underline{\mathbf{Q}}_{\text{coflu}}(\theta)=N\mathbf{Q}_{\text{coflu}}(\theta)$, where $\mathbf{Q}_{\text{coflu}}(\theta)$ is the single-sample quantum Fisher information matrix with the elements
@@
$\tilde{\underline{\rho}}_{P}^{\text{blink}}=\tilde{\rho}_{P1}^{\otimes N_{1}}\otimes...\otimes\tilde{\rho}_{PK}^{\otimes N_{K}}$ (64)
and the quantum Fisher information matrix is $\underline{\mathbf{Q}}_{\text{blink}}(\theta)=\sum_{k}N_{k}\mathbf{Q}^{[k]}(\theta)$, where $\mathbf{Q}^{[k]}(\theta)$ is the single-sample quantum Fisher information matrix in the $k$th window, where just the $k$th source is active:
@@
For $\underline{\mathbf{Q}}_{\text{coflu}}(\theta)$ defined with the object state (62) and $\underline{\mathbf{Q}}_{\text{blink}}(\theta)$ defined with the object state (64), the following inequalities hold:
@@
    $\underline{\mathbf{Q}}_{\text{blink}}(\theta)>\underline{\mathbf{Q}}_{\text{coflu}}(\theta).$ (66)
@@
    $\operatorname{Tr}\{\underline{\mathbf{Q}}_{\text{blink}}(\theta)\}>\operatorname{Tr}\{\underline{\mathbf{Q}}_{\text{coflu}}(\theta)\}.$ (67)
@@
As a corollary to Theorem 4, we have $\underline{\mathbf{Q}}_{\text{blink}}^{-1}(\theta)<\underline{\mathbf{Q}}_{\text{coflu}}^{-1}(\theta)$ in case A. Does it mean that the localization efficiency is always higher in the blinking scenario for any choice of measurement? In the case of Fisher information discussed in Sec. IV, the answer was positive, because the Cramér-Rao bound is always asymptotically saturable. The quantum Cramér-Rao bound, however, is asymptotically saturable only in the case where the equality
$\operatorname{Tr}\left\{\left(\mathcal{L}_{i}\mathcal{L}_{j}-\mathcal{L}_{j}\mathcal{L}_{i}\right)\tilde{\underline{\rho}}_{P}(\theta)\right\}=0$ (68)
@@
$|\Psi_{k}\rangle=\iint\limits_{-\infty}^{+\infty}dxdy\psi(x-x_{k},y-y_{k})a^{\dagger}(x,y)|\text{vac}\rangle.$ (69)
@@
where $i$ takes the values $x_{k},y_{k}$ and the states $|\partial_{x}\Psi_{k}\rangle$ and $|\partial_{y}\Psi_{k}\rangle$ are defined similarly to Eq. (69) but with replacements $\psi(x,y)\to\partial_{x}\psi(x,y)$ and $\psi(x,y)\to\partial_{y}\psi(x,y)$, respectively. Note, that the square of the function $\psi(x,y)$ is normalized to unity, but the square of its derivative is not necessarily that, i.e., the state $|\partial_{i}\Psi_{k}\rangle$ is not normalized in general. Differentiating the norm $\langle\Psi_{k}|\Psi_{k}\rangle=1$, we obtain $\operatorname{Re}\left\{\langle\partial_{i}\Psi_{k}|\Psi_{k}\rangle\right\}=0$, and since the impulse-response function is assumed to be real, $|\partial_{i}\Psi_{k}\rangle$ is orthogonal to $|\Psi_{k}\rangle$. Substituting Eq. (70) into Eq. (68), we ascertain that the commutator average is zero. It means that $x_{k}$ and $y_{k}$ are compatible for a simultaneous quantum estimation in the window where their source is active. The same argument applies to the locations of all other sources.
[…]
It is easy to see that the property of quantum translation and rotation invariance, formulated in Definition 7, is reached for a rotationally invariant impulse-response function, $\psi(\mathbf{R}r)=\psi(r)$, where $\mathbf{R}$ is a matrix of rotation on the $(x,y)$ plane. Indeed, from Eqs. (69) and (70), we obtain $\mathcal{Q}_{2}^{(\mathcal{S})}\left[\tilde{\rho}_{Pk}\right]_{ij}=J_{ij}$, where